Mars has a radius of 3,390 km and a mass of 6.39 x 1023 kg.

a) What is the gravitational field strength on Mars?

b) How much would a 80kg person weigh on Mars?

 

 

 

 

 

 

 

 



1
Expert's answer
2021-06-09T08:04:55-0400

a)

"g=G\frac{M}{R^2},"

"g=6.67\cdot 10^{-11}\cdot \frac{6.39\cdot 10^{23}}{3.39^2\cdot 10^{12}}=3.71~\frac{N}{kg},"

b)

"P=mg,"

"P=3.71\cdot 80\approx300~N."